Horizons has a manufacturing plant located 300 mi from company headquarters. Their corporate pilot must fly from headquarters to the plant and back
in 4 hr. If there is a 20-mph headwind going and a 20-mph tailwind returning, how fast must the plane be able to travel in still air?

Answer by richwmiller(9144)

x+y=4
(p+20)*x=300
(p-20)*y=300
p=152.621, x=1.73791, y=2.26209